### 21 Feb 2024 v1

# Kuaiji: the First Chinese Accounting Large Language Model

Abstract
Large Language Models (LLMs) like ChatGPT and GPT-4 have demonstrated impressive proficiency in comprehending and generating natural language. However, they encounter difficulties when tasked with adapting to specialized domains such as accounting. To address this challenge, we introduce Kuaiji, a tailored Accounting Large Language Model. Kuaiji is meticulously fine-tuned using the Baichuan framework, which encompasses continuous pre-training and supervised fine-tuning processes. Supported by CAtAcctQA, a dataset containing 70,000 genuine accountant-client dialogues, Kuaiji exhibits exceptional accuracy and response speed. Our contributions encompass the creation of the first Chinese accounting dataset, the establishment of Kuaiji as a leading open-source Chinese accounting LLM, and the validation of its efficacy through real-world accounting scenarios.
1 Introduction
Large Language Models (LLMs), such as ChatGPT OpenAI (2022) and GPT-4 OpenAI (2023), have made remarkable strides in pivotal areas. Through extensive pre-training on massive text corpora, they’ve demonstrated exceptional performance across various downstream tasks Kaplan et al. (2020), highlighting their vast potential in understanding and generating natural language Vu et al. (2024). However, despite their progress, LLMs face significant challenges, including the need to adapt to specific tasks and domains Ziegler et al. (2020); Wang et al. (2023), the risk of factual inaccuracies (hallucinations) He et al. (2022), and limitations in handling highly specialized queries Kandpal et al. (2023). These challenges compromise their reliability, especially in fields where accountability and trustworthiness are paramount, such as biomedicine Ji et al. (2023); Song et al. (2024), law Cui et al. (2023a, b), and finance Zhang et al. (2023a); Yang et al. (2023b). Nonetheless, there remains a gap in LLMs tailored for accounting professionals. While many major companies are working on this, most of these models remain closed-source, resulting in a lack of publicly available specialized models for the accounting domain and limiting the widespread application of LLMs in this field.
As in accounting field, people use accounting to track their financial behavior and help them with financial decisions Penman (2010). It’s common sense that accounting information is used by investors and managers to track the firm’s performance and value the company. Accounting information disclosure could change the stock price rapidly Zhu and Niu (2016). However, iterations of accounting regulations and vague accounting books have become a problem for people to use accounting information, and understand the relationship of this information with stock price Sunder (1973). In addition, there is a gap between accounting research and practice, although university education can partly narrow the existing gap, a better solution is needed to bridge the gap Almaan et al. (2015).
In this paper, we endeavor to address the challenges previously mentioned by explicitly fine-tuning a specialized Accounting Large Language Model, named Kuaiji. Based on the Baichuan framework, Kuaiji covers the entire pipeline from continuous pre-training to supervised fine-tuning. Accompanying this, we curated CAtAcctQA, a dataset comprising 70,000 QA pairs from real accountant-client dialogues across 14 accounting departments. Through three stages of model construction, including continuous pre-training, adaptive learning with domain expert feedback, and optimization, Kuaiji achieved remarkable performance. Evaluated against GPT-4 and human experts, Kuaiji excelled in accuracy and response speed, outperforming other open-source Chinese accounting language models. In summary, our contributions are listed as follows:
  • [leftmargin=*]
  • To the best of our knowledge, we have created the inaugural Chinese accounting dataset, featuring 70,000 real instances spanning 14 distinct accounting subareas.
  • Kuaiji stands as the premier open-source Chinese accounting LLM designed explicitly for accounting tasks, serves as the huge knowledge base and inference tools.
  • Our validation of Kuaiji, conducted through real-world accounting cases, showcases its advantages, as suggested by accounting professionals.
2 Related Work
Large Language Models. The advent of Large Language Models (LLMs), exemplified by breakthroughs like ChatGPT OpenAI (2022) and GPT-4 OpenAI (2023), has ignited a transformative wave in artificial intelligence research. These models have not only revolutionized natural language understanding and generation but have also paved the way for groundbreaking advancements across diverse applications. Despite the proprietary nature of some LLMs’ training methodologies, the emergence of open-source alternatives such as Baichuan Yang et al. (2023a), LLaMA Touvron et al. (2023), Bloom Workshop et al. (2023), and Falcon Penedo et al. (2023) has democratized access to state-of-the-art language modeling capabilities. Generative LLMs, or Language Models, exhibit remarkable capabilities in generating text that is both coherent and contextually relevant. Through pretraining on extensive text corpora and fine-tuning for alignment with human instructions, they excel in producing human-like text in response to given prompts or inputs. In essence, LLMs operate by modeling the probability of a sentence $s$ (comprising a sequence of word tokens $q_{1},q_{2},\ldots,q_{n}$) as $p(s)=\prod_{i}^{n}p(q_{i}\mid q_{<i})$. Here, $q_{i}$ represents the $i$-th token of the sentence $s$, while $q_{<i}$ denotes the partial word token sequence preceding the $i$-th step.
Finetune. The most direct approach to injecting domain-specific knowledge into LLM is to finetune these models using high-quality datasets, enabling them to possess domain-specific QA capabilities through supervised learning. HuatuoGPT Zhang et al. (2023b) leverages a mixture of medical knowledge generated by ChatGPT and real-world medical consultation data to achieve performance surpassing that of ChatGPT in the field of Chinese medical knowledge QA. ChatDoctor Li et al. (2023) finetunes LLaMA with 100,000 real-world data samples from online medical consultation platforms and employs a web engine search to assist in generating responses. KALA Kang et al. (2022) incorporates domain-specific knowledge in the form of triplets from KGss into the finetuning process, mitigating training expenses by freezing certain parameters. It is worth noting, however, that the acquisition and balanced curation of high-quality datasets during the finetuning process can be quite challenging, which directly impacts the performance of the finetuned LLMs Zhang et al. (2023b). Besides, substantial computational resources are required for fine-tuning base models, presenting challenges for frequent knowledge updates.
3 Dataset Construction
In this section, we present an exposition on the datasets utilized for training the Kuaiji model. To enhance the efficacy of model training, we meticulously curated two distinct datasets tailored for various stages of training: a continuous pre-training dataset and a supervised fine-tuning dataset. These datasets encompass fundamental and specialized knowledge spanning disciplines such as accounting, auditing, financial management, economics, and economic law. Rigorous manual curation and validation were employed to ensure the quality of these datasets, resulting in the procurement of a high-quality corpus.
3.1 Construction of Continuous Pre-training Dataset
The accounting domain presents a demanding landscape characterized by its need for specialized expertise and its expansive scope covering diverse topics. Consequently, the construction of datasets for model development assumes paramount importance, necessitating a meticulous curation process. In our endeavor to meet the training requisites of our model, we placed particular emphasis on sourcing materials from authoritative sources within academia. Specifically, we selected university textbooks and academic papers from disciplines such as accounting, auditing, tax law, and economic law to form the foundation of our Continuous Pre-training dataset. Additionally, we incorporated supplementary materials from economics, business management, and corporate strategy to enrich the dataset further and ensure its professional integrity.
Diversification of the dataset was achieved through comprehensive data collection efforts spanning various channels. These channels included specialized financial investigation reports, annual reports of publicly listed companies, national statistical bureaus, financial commentaries from columns, media coverage of financial news, and other open datasets. The diverse array of data sources not only enhances the dataset’s breadth but also imbues it with real-time relevance, mirroring the complexities of real-world scenarios. Moreover, these sources serve as complementary resources to traditional academic materials, fostering a comprehensive understanding of professional knowledge within the model.
In tandem with efforts to enhance dataset diversity, rigorous data cleansing procedures were implemented to ensure the dataset’s quality and integrity. This involved meticulous error correction, noise reduction, and consistency checks. Additionally, stringent measures were taken to safeguard data privacy, including the removal of sensitive information and anonymization techniques. Furthermore, expert validation from professionals spanning accounting, economics, law, and related fields was sought to provide an additional layer of scrutiny, thereby upholding the dataset’s professionalism and accuracy. The types and sizes of the datasets we finally obtained are presented in Table 1.
Dataset Type Size
University Textbook
Specialized 324MB
Academic Paper
Specialized 206MB
Accounting wiki
wiki Data 78MB
Company Financial Statements
Related field 173MB
National Statistical Data
Related field 7MB
Financial Survey Report
Related field 16MB
Financial News and Commentary
Related field 24MB
Table 1: KuaiJi pre-training data statistics and sources.
3.2 Construction of Supervised Fine-Tuning Dataset
Fine-tuning is crucial for the training of large models, as it enables models to better understand and generate text relevant to specific domains or tasks by training on domain-specific data. However, due to the scarcity of fine-tuning datasets in certain domains, we opted to utilize domain-specific practice exercises and exam papers to aid in dataset construction. Nevertheless, the existing pool of practice exercises may not cover all necessary aspects comprehensively. To address this shortfall, we devised a specialized prompt to assist in generating the fine-tuning dataset with the help of GPT.
3.2.1 From Test
In the selection of our original dataset, we included practice questions from various sources, including the Certified Practising Accountant (CPA) exam for junior, intermediate, and senior levels, graduate entrance exams for accounting and management fields, as well as exercise collections from professional textbooks and industry-standard problem compilations. To ensure the accuracy of dataset instructions and input-output pairs, we excluded ambiguous questions that might lead to ambiguous outputs during the dataset construction process. Instead, we focused on selecting questions with clear and definitive answers. Moreover, in our design process, we aimed to cover various aspects and difficulty levels relevant to the model’s learning requirements and target tasks comprehensively.
Similarly, expert validation was conducted to ensure the clarity and conciseness of the questions, minimizing ambiguity and ensuring the explicit understandability of the tested content. We prepared a total of 15,677 data points derived from these practice question collections.
3.2.2 From GPT
We selected our self-constructed Continuous Pre-training dataset and supplemented it with fine-tuning data generated by GPT. Initially, we designed a prompt (as depicted in Table 2) to automatically generate instructions, input, and output by analyzing the input text and formatting the output accordingly. Additionally, we implemented a sliding window mechanism to ensure that GPT reads the text with a fixed window size (in this case, the number of lines) and slides at an appropriate distance. This approach guarantees that the generated content is contextually relevant, thereby enhancing logical coherence, consistency, and accuracy while minimizing the risk of semantic discontinuity. Through comparative testing, we determined that the optimal window size and sliding stride are 50 and 30, respectively.
Table 2: Instruction Construction Prompt.
[Task Description]
You are currently an expert in the field of accounting and finance and have been tasked with providing 10 diverse task instructions based on [Background Knowledge]. The following [Requirements] outline the criteria you need to meet when providing instructions, as detailed in the [Reference Format].
[Requirements]
1) Avoid repeating verbs in each instruction and strive for diversity in both instructions and tone.
2) Instructions should encompass various types of tasks, including brainstorming, open and closed QA, rewriting, extraction, generation, classification, chat, and summarization.
3) Instructions should be in Chinese and consist of 1 to 2 sentences. Imperative or interrogative sentences are allowed, without line breaks.
4) You should generate appropriate input for the instruction, including specific examples that involve real data to make the instruction challenging.
5) Not all instructions require input. For example, for common sense information inquiries, such as “what is the highest mountain in the world,” simply state “No Input” in the input field. Text materials (e.g., articles, links) should provide examples directly in the input section. Other media types, like audio, images, videos, or links, do not meet the requirements.
6) The output should be an appropriate response to the instruction and input.
[Reference Format]
[ “Instruction”: “According to the provided accounting rules, classify the following words into the correct categories: Assets, Equity, Expenses, Revenue.”, “Input”: “Understanding the accounting rules: Debit increases and credit decreases assets and equity oppositely. Fixed assets are the same overall, remember details clearly and don’t confused. Distinguish between profit and loss accounts; expenses and revenue are not the same.”, “Output”: “Assets: Fixed assets Equity: Debit increases and credit decreases Expenses: Expenses Revenue: Revenue” ,
{ “Instruction”: “Based on the following text, summarize the important characteristics or definitions of the relationship between tax law and tax legislation.”, “Input”: “Main question: Is national taxation based on political power or property rights? Subquestion 1: Expression of tax law and tax legislation relationship.”, “Output”: “Tax law fundamentally embodies a distribution relationship. The content of tax legislation relationship is the most substantial aspect of tax legislation and the essence of tax law.” }]
[Background Knowledge]
Financial statement auditing is the core business of certified public accountants. A financial statement audit refers to the provision of reasonable assurance by a certified public accountant on whether the financial statements are free from material misstatement, actively expressing opinions, and enhancing the level of trust of expected users other than management in the financial statements…
4 Methodology
This section delves into the fine-tuning process of Kuaiji, spanning two key stages: Continuous Pre-Training (CPT) in Section 4.1 and Supervised Fine-Tuning (SFT)in Section 4.2. Each stage is discussed sequentially to mirror the research workflow. Additionally, we provide a comprehensive method flowchart, depicted in Figure 1, to offer a visual representation of the methodology employed in Kuaiji.
Figure 1: The overall flowchart of constructing Kuaiji.
4.1 Continuous Pre-training
The high-quality pre-training corpus can greatly improve the performance of LLM and even break the scaling laws to some extent Gunasekar et al. (2023). Among them, continuous pre-training is a crucial phase where the Kuaiji model undergoes extensive training on vast and diverse unlabeled datasets. This process spans multiple iterations, each aimed at refining the model’s language understanding capabilities. Initially, the Kuaiji is initialized with the pre-trained weights of basic LLMs and learns to predict missing words or segments within sentences using self-supervised learning objectives such as masked language modeling (MLM) and next-sentence prediction (NSP). Through exposure to a wide variety of textual data sources, as suggested in Section 4.1, the model gradually acquires a rich domain understanding of language structure, semantics, and context in the accounting domain. Additionally, techniques like attention mechanisms and multi-layer architectures are employed to capture complex linguistic patterns and dependencies.
4.2 Supervised Fine-Tuning
Supervised fine-tuning is a pivotal step in refining the Kuaiji model for specific accounting tasks. Leveraging task-specific labeled datasets from both human and strong LLMs, such as the meticulously curated CAtAcctQA dataset and the distilled data from GPT, this phase focuses on adapting the pre-trained model to the intricacies of accounting language and domain-specific conventions. With high-quality domain dialogue data, the model can effectively invoke the accounting knowledge accumulated during pre-training, thereby understanding and responding to users’ queries. SFT involves adjusting the model’s parameters through gradient-based optimization techniques, guided by the labeled data’s ground truth annotations. By iteratively feeding batches of labeled examples to the model and updating its weights through back-propagation, the model learns to refine its internal representations to better align with the target accounting tasks. This process allows Kuaiji to capture nuances in accounting terminology, understand contextual cues unique to financial documents, and effectively handle domain-specific queries and tasks. The supervised fine-tuning phase is instrumental in enhancing Kuaiji’s accuracy, responsiveness, and overall performance in accounting language processing, making it well-suited for real-world applications in finance and accounting domains.
Moreover, to avoid the collapse of inherent capabilities rather than learning substantive ones Gudibande et al. (2023); Shumailov et al. (2023), we mix the domain-specific instructions with general ones, with the proportion of 80%:20% Chen et al. (2023); Dong et al. (2024).
4.3 Training Strategy: QLoRA
During model training, we adopt QLoRA due to its ability to mitigate memory limitations and performance degradation encountered during CPT and SFT phases. By leveraging low-precision quantization and optimized memory management, QLoRA effectively minimizes memory usage while preserving model accuracy. This empowers us to efficiently fine-tune models even on hardware-constrained devices, facilitating the deployment and utilization of large-scale models in practical applications.
The 4-bit NormalFloat quantization technique introduced in QLoRA Dettmers et al. (2023) is based on Quantile Quantization and Block-wise Quantization. It involves mapping the parameters of the neural network to a 4-bit representation while preserving the distribution of the original parameters. This is achieved by dividing the parameter space into quantiles and mapping each quantile to a specific value.
Using the components described above, QLoRA for a single linear layer in the quantized base model with a single LoRA adapter Hu et al. (2021) is defined as follows:
$Y^{\text{BF16}}=X^{\text{BF16}}\cdot\text{doubleDequant}(c^{\text{FP32}},c^{k\text{-bit}},W^{\text{NF4}})+$ (1)
$X^{\text{BF16}}\cdot L^{\text{BF16}}_{1}\cdot L^{\text{BF16}}_{2}$
Where $\text{doubleDequant}(\cdot)$ is defined as:
$\text{doubleDequant}(c^{\text{FP32}},c^{k\text{-bit}},W^{k\text{-bit}})=\text{dequant}\bigl{(}\text{dequant}$ (2)
$(c^{\text{FP32}},c^{k\text{-bit}}),W^{\text{4bit}}\bigl{)}=W^{\text{BF16}}.$
5 Experiments
5.1 Training Details
Our model is based on Baichuan2-13B-chat, a specialized Chinese Large Language Model (LLM) consisting of 13 billion parameters, specifically designed for conversational tasks. The training process utilizes 4 A100-80G GPUs with parallelization techniques within the Baichuan2 framework. During training, we implement the low-rank adaptation (LORA) parameter-efficient tuning method with rank equal to 8 to optimize model performance. This methodology is facilitated through the Transformers4 and PEFT5 libraries. To ensure efficient resource utilization, we adopt fp16 precision with ZeRO-2, along with gradient accumulation strategies. Additionally, to manage computational complexity, we limit the length of individual responses to 4096 tokens. Optimization techniques include the AdamW optimizer, a dropout rate of 0.1, and a cosine learning rate scheduler. A portion of the training dataset (10%) is reserved for validation, and the best-performing checkpoints are selected as the final model. Mechanisms are also in place to handle loss during gradient explosions and to adjust the learning rate for training stability. The hyperparameters for each stage of training are iteratively refined to ensure convergence within an effective loss range.
5.2 Baselines
To conduct a thorough evaluation of our model, we have chosen a variety of LLMs with different parameter scales as baselines for comparison. These include both general-purpose and account-specific LLMs:
  • [leftmargin=*]
  • ChatGPT OpenAI (2022): This widely recognized LLM boasts approximately 175 billion parameters. While not explicitly trained for medical applications, it has demonstrated remarkable performance in account dialogue tasks.
  • Baichuan2-13B Yang et al. (2023a): We also include the base Baichuan2-13B for comparison.
6 Conclusion and Future Work
In summary, although Large Language Models (LLMs) such as ChatGPT and GPT-4 have showcased remarkable capabilities in natural language tasks, their adaptation to specialized fields like accounting poses significant challenges. In response, we have introduced Kuaiji, a finely-tuned Accounting Large Language Model developed within the Baichuan framework. Through continuous pre-training and supervised fine-tuning, Kuaiji has exhibited outstanding accuracy and responsiveness, underpinned by the CAtAcctQA dataset. Our endeavors have led to the creation of the inaugural Chinese accounting dataset and the establishment of Kuaiji as a premier open-source Chinese accounting LLM, validated through real-world accounting scenarios. Looking forward, our future endeavors will concentrate on integrating Reinforcement Learning from Human Feedback (RLHF) to enhance model refinement. Moreover, efforts will be dedicated to expanding the dataset to encompass a broader and more diverse range of data, ensuring Kuaiji’s ongoing enhancement and effectiveness in processing accounting language across diverse applications.